OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 729 m3?

OpenStudy (anonymous):

@ilovehorses14

OpenStudy (anonymous):

@jabez177

OpenStudy (ilovehorses14):

ok hold on a sec

OpenStudy (ilovehorses14):

V = x^3 = 729 x^3 = 9^3 x = 9 m

OpenStudy (anonymous):

thank you

OpenStudy (ilovehorses14):

yw

OpenStudy (ilovehorses14):

:3